OpenStudy (anonymous):

if sin theta = 1/3 what is cos(90 - theta)

OpenStudy (lalaly):

\[Cos ( 90° - \theta) = \sin \theta =1/3\]

OpenStudy (anonymous):

can u explain a little XD

OpenStudy (anonymous):

she explained it!

OpenStudy (lalaly):

yes i can explain it jus a min

OpenStudy (anonymous):

thank you :)

OpenStudy (lalaly):

Draw a right triangle. Let one of the acute angles be x Then the other acute angle must be (90-x), since the sum of the degrees in the triangle must be 180.